What score is needed to be in the top 10% of a distribution with a mean = 500 and standard deviation = 100?
Proportion
Draw

Answer by stanbon(75887)

What score is needed to be in the top 10% of a distribution with a mean = 500 and standard deviation = 100?
---
Using x = zs+u
x = 1.2816*100 + 500 = 128.16+500 = 628.16
================
Cheers,
Stan H.
================